# CroCoDai: A Stablecoin for Cross-Chain Commerce

### v1

@@ 5.1 Safety Analysis
so that $D^{-}_{t}$ represents the debt associated with affected CDPs and $D^{+}_{t}$ the debt of unaffected CDPs at time $t$. We consider an attack that occurs at time $1$, so that $C^{-}_{1}=0$, $C^{+}_{1}=C^{+}_{0}$, $D^{-}_{1}=D^{-}_{0}$, and $D^{+}_{1}=D^{+}_{0}$. We can then prove the following theorem (see Appendix D):
@@
Individual Compromised Coin Chain. If a majority of consensus power on a coin chain $b^{\prime}\in[B]$ is compromised then the consensus nodes could perform a 51% attack in the following way: send some amount of stablecoins to another chain, wait for the transaction to be committed by the relay network, then fork the compromised coin chain and undo the outgoing transfer request and the burning of the blockchain tokens. Similarly, they could undo the funding of a CDP. As a consequence, stablecoins are created on another chain without backing collateral. In the worst case, this attack would be performed simultaneously on all of the token types supported by the compromised chain $b^{\prime}$. Let $M^{\prime}=\{m\in[M]:b_{m}=b^{\prime}\}$ be the set of affected tokens. We again consider a worst-case attack that has been completed at $t=1$. Let $h^{\prime}$ be the total amount of stablecoins created in the attack, so that $D^{-}_{1}=D^{-}_{0}+h^{\prime}.$ We can then prove the following theorem (see Appendix D):
@@
Compromised Price Feeds. Finally, we consider an attack in which a dishonest minority of price oracles tries to corrupt the price maintained by a vault. This could cause CDPs to be liquidated even if they are not underwater, thus violating safety of collateral.
Let $p$ be the base price for token $m\in[M]$ at time $t\geq 0$ (we omit $m$ and $t$ from the notation in this section as they are not relevant to the analysis). Let $p_{o}$ be the price returned by oracle $o\in[O]$, so that $p_{o}=p+\Delta_{o}$ as discussed in Section 3.2. In the following, we assume some subset $O^{\prime}\subset[O]$ is controlled by malicious parties who aim to convince a vault in CroCoDai that of a price $p^{\prime}$ that is different from the base price $p$. Let $p^{*}$ be the actual median price among all price feeds, including those that have been distorted. We can then prove the following theorem (see Appendix D):
@@
5.2 Financial Analysis: Approach
@@
In this section, we investigate the probability that CroCoDai remains fully backed even in the presence of standard price fluctuations. In particular, we will estimate this probability for several choices of supported collateral tokens: 1) a single token type, 2) tokens that exist on the Ethereum blockchain, and 3) tokens that exist on a variety of independent blockchains. The second probability corresponds to the safety of the single-chain DSS, whereas the third option corresponds to the safety of CroCoDai. The purpose of this section is to demonstrate that the third option has considerably lower risk than the second, using multi-chain collateral baskets that are similar to those observed for the DSS in practice (see Table 8 in Appendix E). We discuss our approach in Section 5.2 and present our empirical results in Section 5.3.
As is common in financial analysis, we will refer to a basket of collateral tokens as a portfolio, and to the relative increase in value of a token type over one time slot as a return. We will also discuss a method based on convex optimization that allows protocol engineers to set the system parameters such that the risk of system failure is minimized.
@@
In our simulation, we represent the portfolio as relative proportions of different assets. We define a vector $\vec{v}$ that denotes the value of individual assets in the portfolio as $\textbf{v}=(v_{1},v_{2},\ldots,v_{M})^{T},$ where $v_{i}$ is value of asset $i$. Similarly, we define a vector of random variables for log returns of individual assets $\textbf{X}=(X_{1},X_{2},\ldots,X_{M})^{\intercal},$ where $X_{m}$ is the random variable for log returns of asset $m$. We then simulate the price of each asset as an individual GBM with Student’s $t$-distribution for the return after each time slot, as discussed in Section 3.2. To accurately simulate price fluctuations, we must consider the correlation between the prices – to do so, we use Cholesky decomposition to compute a matrix $\mathbf{L}$ that satisfies $\textnormal{Cov}(\mathbf{X})=\mathbf{L}\mathbf{L}^{T}$. A matrix of random samples $\mathbf{R}$ can then be generated as $\mathbf{R}=\mathbf{LT}$, where the matrix $\mathbf{T}$ contains a $t$-distributed random sample for each asset type and simulated time slot. In Appendix B we discuss alternative simulation methods based purely on historical data and the standard GBM.
Choosing the System Parameters. In CroCoDai, the values $\alpha_{m}$, $m\in[M]$, which represent the fraction of each asset $m$ in the total portfolio, are not set by the protocol engineers – instead, they emerge naturally as users create CDPs and withdraw stablecoins using the currencies of their choosing. However, protocol engineers can influence the range of potential portfolios by adjusting the debt ceiling $\zeta_{m}$ for each token $m$.
When optimizing the portfolio of cryptocurrencies for a stablecoin, the protocol engineers’ main goal should be to minimize the portfolio’s volatility. In principle, a positive mean drift in the returns is beneficial, as that would imply that the over-collateralization grows naturally over time. However, the mean return over our observation period was close to zero, and for risk analysis, with cross-chain commerce as our driving use case, we focus mainly on short-term price changes. In this context, the mean return is not a meaningful optimization criterion, so we focus on minimizing the volatility instead.
Assuming a normal or $t$-distribution of returns, we can calculate the covariance matrix $\text{Cov}(\mathbf{X})$ as discussed previously. We can then use convex optimization to find a portfolio that minimizes total volatility. We can write this optimization problem as
$\begin{array}[]{rcl}\displaystyle\min_{\textbf{v}}&\textbf{v}^{T}\text{Cov}(\mathbf{X})\textbf{v},&\\[5.69046pt] \text{s.t.}&\displaystyle\sum_{i=1}^{M}v_{i}=1,&\\[5.69046pt] &\displaystyle 0\leq v_{i}\leq\lambda_{i}&\forall i\in[M],\end{array}$ (4)
where we consider $M$ different cryptocurrencies and $v_{i}$ denotes the fraction of the portfolio invested in asset $i\in[M]$. The bounds $0\leq\lambda_{i}\leq 1$ enforce maximum contributions for the different individual asset types, to avoid portfolios with a small number of low-volatility tokens (such portfolios would be vulnerable to black-swan events as discussed in Section 5.1). Given a solution $v_{1},\ldots,v_{M}$ to (4), we can enforce the optimal portfolio up to some margin by setting $\zeta_{m}=(1+\beta)v_{m}$ where $\beta>0$ is the flexibility margin: a higher margin encourages more users to create CDPs, but also allows for larger deviations from the optimal portfolio.
5.3 Financial Analysis: Evaluation
@@
Description of Data. We have collected data using the CoinMarketCap API over six different periods that span a total of 1264 hours of data, separated into 5-minute slots. The duration of these periods can be found in Table 9 in Appendix F. The final period includes the collapse of the FTX exchange on November 11, 2022. In each time slot, we collected price information of 100 tokens, namely those that were the 100 largest in terms of market capitalization during the first time slot. A complete list of the included token symbols, and the observed correlation between them, can be found in Figure 6 in Appendix F.
@@
  1. 1.
    D-All, which approximates the collateral portfolio of the DSS as of early 2023 (see Table 8 in Appendix E). (We replace GUSD, which was not in our dataset, with another fiat-backed stablecoin (USDT), and Maker’s real-world assets with gold-backed PAXG.)
  2. 2.
@@
The exact compositions of these portfolios can be found in Table 10 in Appendix G.
In our experiments, the expected over-collateralization to avert system failure is $\theta=1.1$. For each table entry, we calculate the probability of failure within 1 day, i.e., 288 5-minute time slots, using 100 000 simulation runs. In each run, we simulate the evolution of token prices by randomly drawing, in each consecutive time slot, the price change for each token type in the portfolio. We then compute the risk of failure as the fraction of runs that resulted in failure among the total number of runs. The results are displayed in Table 1Appendix B includes similar tables with different simulation methods and confidence interval widths.
Table 1: Chance of failure for single asset portfolios and initial over-collateralization.
@@
From Table 1, we observe a considerable difference between the failure probabilities of the single-token portfolios. The most high-risk portfolio was Solana (SOL), which was severely affected by the collapse of FTX, since it was one of FTX’s most prominent backers [11]. Even for an initial collateralization $\gamma^{\prime}=1.5$, a pure SOL portfolio would have a failure probability of more than $4\%$ after a day, given the data from our observation period. In contrast, BTC and TRX were the most stable cryptocurrencies over this period.
Figure 3: Correlation between the different token types.
@@
Furthermore, we observe from Table 1 that the DSS’s approximate portfolio has a very low probability of failure due to $80\%$ of its collateral being stablecoins and other pegged assets. In fact, for $\gamma^{\prime}=1.2$, its observed failure probability (0.001 for D-All) was roughly 40-50 times lower than for the pure crypto-backed portfolios (0.051, 0.040, and 0.050 respectively for D-Erc, D-Mix1, and D-Mix2). For the pure crypto-backed portfolios, we find that D-ERC has slightly lower failure probabilities than a pure ETH portfolio, which can be explained by the higher share of low-volatility WBTC in the portfolio. Similarly, C-Mix1 and C-Mix2 portfolios have lower risk than the D-ERC due to the inclusion of more WBTC. We do not observe much benefit from the inclusion of more different asset types due to the high correlation between the included tokens. However, when comparing the C-Mix1 and C-Mix2 portfolios we do note that C-Mix2 has a lower failure probability despite the inclusion of less low-volatility WBTC. The reason is the inclusion of TRX, which is not only similarly low-risk as WBTC, but which also has a relatively low correlation with the other asset types – this can be seen in Figure 2(a). We also note that the relative differences become starker when $\gamma^{\prime}$ increases – e.g., for $\gamma^{\prime}=1.2$, the risk of D-Mix1 is roughly a third higher than C-Mix2 (4.0% vs. 2.9%), whereas for $\gamma^{\prime}=1.3$ the risk is two thirds higher (0.5% vs. 0.3%).
Optimized Portfolios. The portfolios in the previous section approximated the existing DSS portfolio, which is aligned with the willingness of real-world users to deposit collateral of the various types. Although we did observe a slightly lower risk probability for the best multi-chain portfolio (C-Mix2) than for the best single-chain portfolio (D-Mix1), the underlying fundamental issue is how they compare if we were to use the convex optimization method based on (4) to determine low-volatility portfolios.
A first question in understanding this is whether to include stablecoins in the portfolio. If we use convex optimization across the entire dataset of 100 token types, we obtain the portfolio from Table 12 in Appendix G. With the exception of BTT (the BitTorrent token), this portfolio only includes USD- or gold-pegged assets. A crypto-backed stablecoin system that only uses fiat-backed stablecoins as collateral has trivially low price volatility, yet would share most of the drawbacks of a fiat-backed stablecoin as discussed in Section 2 (i.e., higher vulnerability to black swan events that involve off-chain entities). As such, we will also consider a stablecoin that purely uses unpegged cryptocurrency assets as collateral. This represents a worst-case scenario in terms of volatility risk: in practice, an appropriate mixture of stablecoins and unpegged cryptocurrencies can be chosen. If we restrict the portfolios to the same tokens as the portfolios of Table 10, we obtain the optimal portfolios D-Opt and C-Opt of Table 11 (both in Appendix G). We also include the portfolio A-Opt which may contain any of the Top 100 tokens except those already pegged to a real-world asset. Since we observe that the optimal portfolios tend to restrict themselves to a handful of tokens (e.g., only WBTC and GNO/TRX for D-Opt and C-Opt), we also set a threshold of $\lambda=0.2$ for D-Opt and C-Opt and $0.1$ for A-Opt to avoid high exposure to events on a single blockchain.
Table 2: Chance of failure for given portfolio and initial over-collateralization.
D-Opt C-Opt A-Opt risk vs. D-Opt
$\gamma^{\prime}$ $\lambda=0.2$ $\lambda=1$ $\lambda=0.2$ $\lambda=1$ $\lambda=0.1$ $\lambda=0.2$ $\lambda=1$ C-Opt A-Opt
1.2 0.03419 0.01423 0.02103 0.00934 0.00895 0.00624 0.00617 -38.49% -81.75%
1.3 0.00397 0.00133 0.00232 0.00131 0.00076 0.00055 0.00081 -41.56% -86.14%
1.4 0.00115 0.00058 0.00087 0.00056 0.00030 0.00020 0.00024 -24.34% -82.61%
1.5 0.00057 0.00027 0.00037 0.00024 0.00014 0.00008 0.00012 -35.08% -85.97%
The results of our optimal portfolio simulations are displayed in Table 2. The final (bold) column displays the decrease in terms of failure risk estimates for C-Opt and A-Opt versus D-Opt for $\lambda=0.2$. In general, we find that the optimized portfolios have slightly lower failure probabilities than the DSS-based ones from the previous section (e.g., for $\gamma^{\prime}=1.1$ and $\lambda=0.2$, we obtain 0.034 for D-Opt vs. 0.040 for D-Mix1, and 0.021 for C-Opt vs. 0.029 for C-Mix2). The C-Opt portfolios again outperform D-Opt portfolios by roughly a third, confirming our expectation that a multi-chain stablecoin is less vulnerable to price fluctuations than a single-chain one. Furthermore, we find even lower failure probabilities for A-Opt, although this portfolio contains many tokens on blockchains that are not as high-use as the tokens in the other portfolios – e.g., BSV or KLAY vs. BTC and ETH in other portfolios. Less-used tokens may be more likely to experience black swan events as there are fewer incentives to scrutinize smart contract code or the identity of the consensus nodes. In any case, we find that even for pure crypto-backed portfolios (C-Opt) and sufficient over-collateralization ($\gamma^{\prime}=1.4$, which is similar to $\gamma=1.5$ in the DSS and $\theta=1.1$) CroCoDai’s estimated system failure (i.e., de-peg) risk after a day is less than one in a thousand, which may be acceptable for low-to-medium-value commerce (e.g., a ticket auction).
6 Implementation
In this section, we discuss the technical details of three prototype implementations of our stablecoin system. We compare them in terms of the on-chain processing cost (gas) and the off-chain computation time. Both are essential to system performance and thus usability: if the on-chain cost is too high then each transfer will incur large fees, whereas if the computation time is too high then there is a large delay for cross-chain transactions. We focus on the implementation of the relay nodes, because all other components depicted in Figure 2 (the coin, vault, auction, savings, and oracle contracts) can be re-used from the DSS without modification – we only add the relay contracts as wards for the coin contract. As such, we cannot provide a straightforward comparison with the DSS, as the performance of CroCoDai contracts is either exactly the same or there is no DSS equivalent. However, we will discuss the comparison with alternatives from the scientific literature, e.g., zkBridge.
Each of our three designs uses a quorum of trusted nodes, but has a different tradeoff in terms of scalability and generality. Each implementation consists of two smart contracts written in Ethereum’s Solidity language: Relay.sol and CCTransfer.sol, although only Relay.sol differs between implementations. The Relay and CCTransfer contracts interact with an ERC-20 compliant contract for token transfers, Coin.sol, which re-uses the code from the Dai.sol contract in the DSS [41]. Our experiments were run on a Lenovo IdeaPad with an Intel(R) Core(TM) i5-1035G1 CPU (1.00GHz) and 8GB RAM (although the gas costs are machine-independent). The scripts that performed off-chain computations were written in Python and used the Web3 package for most cryptographic operations, while the underlying blockchain was simulated using Truffle’s Ganache tool [21]. For communication about cross-chain transfers, the relay nodes can use a private blockchain, e.g., a private Go-Ethereum blockchain, or a bespoke cross-chain communication framework such as [46].
6.1 Smart Contracts

### v4

@@ 4.3. Sudden Crash of Tokens
so that $D^{-}_{t}$ represents the debt associated with affected CDPs and $D^{+}_{t}$ the debt of unaffected CDPs at time $t$. We consider an attack that occurs at time $1$, so that $C^{-}_{1}=0$, $C^{+}_{1}=C^{+}_{0}$, $D^{-}_{1}=D^{-}_{0}$, and $D^{+}_{1}=D^{+}_{0}$. We can then prove the following theorem (see Appendix A):
@@
4.4. Individual Compromised Coin Chain
If a majority of consensus power on a coin chain is compromised then the consensus nodes could perform a 51% attack in the following way: send some amount of stablecoins to another chain, wait for the transaction to be committed by the relay network, then fork the compromised coin chain and undo the outgoing transfer request and the burning of the blockchain tokens. Similarly, they could undo the funding of a CDP. As a consequence, stablecoins are created on another chain without backing collateral. In the worst case, this attack would be performed simultaneously on all of the token types supported by the compromised chain. Again, let $M^{\prime}\subset M$ denote the set of affected tokens. We again consider a worst-case attack that has been completed at $t=1$. Let $h^{\prime}$ be the total amount of stablecoins created in the attack, so that $D^{-}_{1}=D^{-}_{0}+h^{\prime}.$ We can then prove the following theorem (see Appendix A):
@@ 4.4. Individual Compromised Coin Chain
4.5. Compromised Price Feeds
Finally, we consider an attack in which a dishonest minority of price oracles tries to corrupt the price maintained by a vault. This could cause CDPs to be liquidated even if they are not underwater, thus violating safety of collateral. We note that the DSS already has implemented safety mechanisms for corrupted price oracles, e.g., its Oracle Security Module (OSM). The OSM [^11] https://docs.makerdao.com/smart-contract-modules/oracle-module/oracle-security-module-osm-detailed-documentation increases the available time to detect corrupted price feeds by adding an artificial delay before price information is accepted by the system. Such safety mechanisms are complementary to our analysis in the sense that they make individual price oracles harder to compromise, whereas our analysis shows that attacks against the system as a whole have negligible probability of success if only a minority of oracles is compromised. Furthermore, CroCoDai’s inclusion of global price oracles on minor chains provides another complementary safety mechanism. After all, if price oracles from established global sources are included, then it is harder for attackers to compromise a majority among the full set of oracles than if this set only includes local sources from the minor chain’s less developed DeFi ecosystem.
Let $p$ be the base price for token $m\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}$ at time $t\geq 0$ (we omit $m$ and $t$ from the notation in this section as they are not relevant to the analysis). Let $O$ denote the set of all oracles and $p_{o}$ the price returned by oracle $o\in O$, so that $p_{o}=p+\Delta_{o}$ where $\Delta_{o}$ is a normally distributed random variable as per Section 4.1. In the following, we assume some subset $O^{\prime}\subset O$ is controlled by malicious parties who aim to convince a CroCoDai vault to accept a price $p^{\prime}$ that is different from the base price $p$. Let $p^{*}$ be the actual median price among all price feeds, including those that have been distorted. We can then prove the following theorem (see Appendix A):
@@ 4.5. Compromised Price Feeds
4.6. Comparison to Baselines
In a fully siloed baseline, each coin chain has its own DSS-like stablecoin. If the chain is not compromised, then local and cross-chain transfers are secure (i.e., tokens cannot be stolen or locked permanently), whereas they are not if the chain is compromised. In this sense, the baseline is similar to CroCoDai. The fully siloed system is unaffected by price crashes or platform failures on other blockchains – however, it is less resilient to price crashes and platform failures on its own chain than a system with more diversified collateral: e.g., in CroCoDai, a stablecoin holder on a minor chain that experiences a price collapse would still be able to transfer stablecoins to another chain and exchange them for other tokens. Finally, because CroCoDai supports additional price oracles, it is more difficult for an adversary to compromise a majority of oracles than in the baseline.
The baseline that consists of the DSS with a wider set of wrapped tokens has the same security guarantees as CroCoDai, as it is similarly able to diversify its collateral. However, we will see in Section 6 that this baseline has lower efficiency.
5. Financial Analysis
The main system requirement from Section 3.4, full backing, can be violated not just due to critical attacks and platform failures, but also due to natural price fluctuations in token prices. This vulnerability to price fluctuations is compounded by the strong correlation between the prices of different cryptocurrencies (see Figure 6 in the appendix). In this section, we use a financial risk model to estimate the probability that the system violates full backing over a 1-day period. As is common in financial analysis, we will refer to a basket of collateral tokens as a portfolio, and to the relative increase in value of a token type over one time slot as a return. For our risk analysis, we compare a collateral portfolio that approximates the collateral composition of the existing DSS (without other stablecoins) to a portfolio made up of a wider range of blockchain tokens. In the following, we first present our model for token prices, and then present our risk analysis for the DSS-based portfolios.
5.1. Token Price Model
As mentioned in Section 4, we denote the base price of token $m\in M$ at time $t>0$ by $p_{m\,t}$. To model the evolution of token prices over time, we use Geometric Brownian Motion (GBM). This is a common choice in the financial literature on asset prices (ross2011elementary, ) and related work on stablecoins, e.g., by Salehi et al. (salehi2021red, ), particularly due to the GBM’s role in the Black-Scholes formula for option prices. A stochastic process $S_{t}$ that follows GBM is defined using a stochastic differential equation:
(1) $dS_{t}=\mu S_{t}dt+\sigma S_{t}dW_{t},$
where $\mu$ determines the drift of $S_{t}$ (i.e., its expected change per unit of time), $\sigma$ represents the volatility of $S_{t}$ (i.e., the variance of the change), and $W_{t}$ is the Wiener process that represents standard Brownian motion (ross2011elementary, ). In this model, each token price process $p_{m\,t}$, $t\geq 0$, can be represented using a process defined as in Equation 1. Essentially, this means that the logarithm of the relative difference of token prices, $\log\left(p_{m\,t+t^{\prime}}/p_{m\,t}\right)$, is normally distributed for all $t^{\prime}>0$ with mean and variance dependent on $t^{\prime}$. Unlike (salehi2021red, ), we consider a setting with multiple assets; i.e., we have a stochastic process $S_{m\,t}$ for each token $m$ with its own mean $\mu_{m}$ and variance $\sigma^{2}_{m}$. In this setting, we must also consider the covariance between the token prices (ross2011elementary, ), as cryptocurrency prices are often highly correlated – we will discuss this in more detail below.
The standard GBM as decribed above is based on the normal distribution, i.e., $\log(S_{t+t^{\prime}}/S_{t})$ has a normal distribution. However, it is known from the academic literature that the probability distribution of stock price fluctuations is usually more heavy-tailed and that Student’s $t$-distribution is often a better fit (kelly2014tail, ; li2020review, ). This is also noticeable in the data for price changes in our dataset. As such, we will focus on the $t$-distribution in our analysis. Although the $t$-distribution does not allow for an analytical solution for the asset price change after each individual time slot, we can use Monte Carlo simulation to analyze its behavior, as we discuss in more detail below.
5.2. Risk Analysis: Approach
@@
In this section, we investigate the probability that CroCoDai remains fully backed even in the presence of standard price fluctuations. In particular, we will estimate this probability for several choices of supported collateral tokens: 1) a single token type, 2) tokens that exist on the Ethereum blockchain, and 3) tokens that exist on a variety of independent blockchains. The second probability corresponds to the safety of the single-chain DSS, whereas the third option corresponds to the safety of CroCoDai. The purpose of this section is to demonstrate that the third option has considerably lower risk than the second, using multi-chain collateral baskets that are similar to those observed for the DSS in practice (see Table 5 in Appendix B). We discuss our approach in Section 5.2 and present our empirical results in Section 5.3.
@@
In our simulation, we represent the portfolio as relative proportions of different assets. We define a vector $\vec{v}$ that denotes the value of individual assets in the portfolio as $\textbf{v}=(v_{1},v_{2},\ldots,v_{|M|})^{T},$ where $v_{i}$ is value of asset $i$. Similarly, we define a vector of random variables for log returns of individual assets $\textbf{X}=(X_{1},X_{2},\ldots,X_{|M|})^{\intercal},$ where $X_{m}$ is the random variable for log returns of asset $m\in M$. We then simulate the price of each asset as an individual GBM with Student’s $t$-distribution for the return after each time slot, as discussed in Section 5.1. To accurately simulate price fluctuations, we must consider the correlation between the prices – to do so, we use Cholesky decomposition to compute a matrix $\mathbf{L}$ that satisfies $\textnormal{Cov}(\mathbf{X})=\mathbf{L}\mathbf{L}^{T}$. A matrix of random samples $\mathbf{R}$ can then be generated as $\mathbf{R}=\mathbf{LT}$, where the matrix $\mathbf{T}$ contains a $t$-distributed random sample for each asset type and simulated time slot.
5.3. Risk Analysis: Evaluation
@@
Description of Data. We have collected data using the CoinMarketCap API over six different periods that span a total of 1264 hours of data, separated into 5-minute slots. The duration of these periods can be found in Table 6 in Appendix C. The final period includes the collapse of the FTX exchange on November 11, 2022. Our main motivation was to choose a diverse range of market conditions, including a period of turmoil, to ensure that the estimated system failure probabilities represent a worst-case scenario (i.e., in more typical market conditions, CroCoDai would be safer). In each time slot, we collected price information of 100 tokens, namely those that were the 100 largest in terms of market capitalization during the first time slot. A complete list of the included token symbols, and the observed correlation between them, can be found in Figure 6 in Appendix C.
@@
  1. (1)
    D-All, which approximates the collateral portfolio of the DSS as of early 2023 (see Table 5 in Appendix B). (We replace GUSD, which was not in our dataset, with another fiat-backed stablecoin (USDT), and Maker’s real-world assets with gold-backed PAXG.)
  2. (2)
@@
We use these portfolios to show how the system performs with a mix of different types of collateral: other (fiat-backed) stablecoins, major cryptocurrencies (e.g., ETH/WBTC), and minor ERC-20 tokens. Tokens of the first type typically have much lower volatility, but introduce a dependence on trusted third parties. We will discuss this trade-off in more detail later in this section. The exact compositions of the portfolios can be found in Table 7 in Appendix D.
In our experiments, the expected over-collateralization to avert system failure is $\theta=1.1$. For each table entry, we calculate the probability of failure within 1 day, i.e., 288 5-minute time slots, using 100 000 simulation runs. In each run, we simulate the evolution of token prices by randomly drawing, in each consecutive time slot, the price change for each token type in the portfolio. We then compute the risk of failure as the fraction of runs that resulted in failure among the total number of runs. The results are displayed in Table 1. Since the 95% confidence intervals widths are very small for the results in this table, we have omitted them for brevity.
Table 1. Chance of failure for single-asset portfolios and initial over-collateralization $\gamma^{\prime}$.
@@
Table 2. Chance of failure for single-asset portfolios and initial over-collateralization $\gamma^{\prime}$. In the table, ‘0.0 $\pm$ 0.0’ indicates that system failure events were observed during the 1,000,000 simulation runs, but that the estimated probability has been rounded down to 00, whereas ‘—’ indicates that no failure events were observed.
single-asset C-Mix1, stablecoin contribution $\delta$ $\gamma^{\prime}$ USDT USDC BUSD TUSD DAI $\delta=0.5$ $\delta=0.25$ $\delta=0.1$ $\delta=0.05$ $\delta=0$ vol. 4.85e-5 9.70e-5 4.06e-4 3.48e-4 2.84e-4 7.86e-4 1.17e-3 1.39e-3 1.47e-3 1.55e-3 1.2 0.00001 $\pm$2.2e-6 0.00009 $\pm$9.2e-6 0.00177 $\pm$4.2e-5 0.00103 $\pm$3.2e-5 0.00078 $\pm$2.8e-5 0.00314 $\pm$5.6e-5 0.01198 $\pm$1.1e-4 0.02153 $\pm$1.5e-4 0.02548 $\pm$1.6e-4 0.02976 $\pm$1.7e-4 1.3 0.0 $\pm$0.0 0.00002 $\pm$4.7e-6 0.00041 $\pm$2.0e-5 0.00027 $\pm$1.6e-5 0.00023 $\pm$1.5e-5 0.00047 $\pm$2.2e-5 0.00137 $\pm$3.7e-5 0.00243 $\pm$4.9e-5 0.00304 $\pm$5.5e-5 0.00345 $\pm$5.9e-5 1.4 0.00001 $\pm$3.0e-6 0.00021 $\pm$1.4e-5 0.00012 $\pm$1.6e-5 0.00009 $\pm$9.3e-6 0.00018 $\pm$1.3e-5 0.00052 $\pm$2.3e-5 0.00098 $\pm$3.1e-5 0.00105 $\pm$3.2e-5 0.00119 $\pm$3.5e-5 1.5 0.0 $\pm$0.0 0.00001 $\pm$2.8e-6 0.00012 $\pm$1.1e-5 0.00001 $\pm$8.3e-6 0.00001 $\pm$7.7e-6 0.00009 $\pm$9.6e-6 0.00029 $\pm$1.7e-5 0.00046 $\pm$2.1e-5 0.00033 $\pm$1.8e-5 0.00065 $\pm$2.5e-5
@@
From Table 1, we observe a considerable difference between the failure probabilities of the single-token portfolios. The most high-risk portfolio was Solana (SOL), which was severely affected by the collapse of FTX, since it was one of FTX’s most prominent backers (business_insider, ). Even for an initial collateralization $\gamma^{\prime}=1.5$, a pure SOL portfolio would have a failure probability of more than $4\%$ after a day, given the data from our observation period. In contrast, BTC and TRX were the most stable cryptocurrencies over this period.
(a) Different blockchains
(b) ERC20 tokens
Figure 5. Correlation between the different token types.
@@
Furthermore, we observe from Table 1 that the DSS’s approximate portfolio has a very low probability of failure due to $80\%$ of its collateral being stablecoins and other pegged assets. In fact, for $\gamma^{\prime}=1.2$, its observed failure probability (0.001 for D-All) was roughly 40-50 times lower than for the pure crypto-backed portfolios (0.051, 0.040, and 0.050 respectively for D-Erc, D-Mix1, and D-Mix2). For the pure crypto-backed portfolios, we find that D-ERC has slightly lower failure probabilities than a pure ETH portfolio, which can be explained by the higher share of low-volatility WBTC in the portfolio. Similarly, C-Mix1 and C-Mix2 portfolios have lower risk than the D-ERC due to the inclusion of more WBTC. We do not observe much benefit from the inclusion of more different asset types due to the high correlation between the included tokens. However, when comparing the C-Mix1 and C-Mix2 portfolios we do note that C-Mix2 has a lower failure probability despite the inclusion of less low-volatility WBTC. The reason is the inclusion of TRX, which is not only similarly low-risk as WBTC, but which also has a relatively low correlation with the other asset types – this can be seen in Figure 5a. We also note that the relative differences become starker when $\gamma^{\prime}$ increases – e.g., for $\gamma^{\prime}=1.2$, the risk of D-Mix1 is roughly a third higher than C-Mix2 (4.0% vs. 2.9%), whereas for $\gamma^{\prime}=1.3$ the risk is two thirds higher (0.5% vs. 0.3%).
Comparison to Existing Stablecoins. From Table 1, we observe that D-All has a considerably lower failure risk than the other portfolios. D-All represents the collateral portfolio of the DSS and is the only multi-asset portfolio from Table 1 that includes other stable assets (around 75% as per Table 5). All stablecoins except Dai in our dataset are fiat-backed stablecoins, which exhibit much smaller price volatility than regular tokens, but rely on trusted third parties which expose the system to external black swan events. We conduct experiments to further understand the price volatility. In particular, we estimate the failure risks for single-stablecoin portfolios and for a range of DSS-based portfolios with varying degrees of included stablecoins. For the latter, we consider modifications of the C-Mix1 portfolio, which represents a portfolio similar to Dai’s collateral portfolio but which includes a minor fraction (20%) of tokens from other blockchains. We modify this portfolio by assigning a fraction $\delta$ of fiat-backed stablecoins to it and multiplying the contribution of the other token types by $1-\delta$, for varying values of $\delta$. The low volatility of other stablecoins means that the failure probabilities of single-stablecoin portfolios are very low. Therefore, we conduct 1,000,000 runs for each table entry and display the 95% confidence interval half-width below each estimate in Table 2. We also display the variance, which is strongly related to the failure risks, of each portfolio in the table (under vol.).
From Table 2, we observe that the failure risks of all single-stablecoin portfolios are much lower than those of the single-asset portfolios in Table 1, although there are still considerable differences between individual stablecoins: e.g., BUSD’s volaility is 10 times higher than USDT’s, and, for $\gamma^{\prime}$, BUSD’s estimated 1-day probability of failure is 177 times higher than that of USDT. For the modified C-Mix1 portfolios, we have used USDC as a representative stablecoin because of its high trade volume (there have been over $20$bn USDC in circulation since July 2021) and because its volatility is neither exceptionally high nor low among the included stablecoins. It can be seen that, as expected, the inclusion of other stablecoins leads to low failure risks, although we observe a small decrease for higher values of $\gamma^{\prime}$: in particular, the relative difference between $\delta=0$ and $\delta=0.25$ is roughly 2.5$\times$ (0.02976 vs. 0.01198), but this decreases to 2.24$\times$ for $\gamma^{\prime}=1.5$. For $\delta=0$ and $\delta=0.5$, the relative difference decreases from 9.48$\times$ for $\gamma^{\prime}=1.1$ to 7.22$\times$ for $\gamma^{\prime}=1.5$. Overall, we observe that the inclusion of stablecoin fraction of $25\%$ can reduce the system’s failure risk by more than $60\%$.
We note that a stablecoin system that is overcollateralized through a mixture of fiat-backed stablecoins is more resilient to black-swan events than the individual stablecoins. The reason is that the collapse of any single trusted third party would have a limited impact. However, black swan events may affect multiple fiat-backed stablecoins simultaneously, e.g., due to new legislation affecting stablecoin collateral. Thus, it would still be advisable to ensure that the total contribution of fiat-backed stablecoins is limited. We also note, that, in theory, a principal crypto-backed stablecoin could include other crypto-backed stablecoins as collateral. In this setting, the portfolio of the supporting stablecoins should be included in the portfolio of the principal stablecoin. Finally, protocol designers with substantial fiat assets (e.g., Dai) also have the option to steer collateral portfolios towards those with minimal risk (hajek2024collateral, ), which would reduce the risks associated with the crypto collateral even further.
6. Implementation
In this section, we present a prototype implementation of CroCoDai and compare its performance in terms of both cost and processing time to a baseline that combines the existing DSS with wrapped tokens. We present three different subvariants of the relay chain design with a quorum of nodes (i.e., the offchain variant of Section 3.3). Each subvariant has a different tradeoff in terms of scalability and generality, and consists of two smart contracts written in Ethereum’s Solidity language: Relay and CCTransfer, with only Relay different between the three subvariants. The Relay and CCTransfer contracts interact with an ERC-20 compliant contract for token transfers, Coin, which re-uses code from the Dai contract in the DSS (daisol, ). We discuss the smart contracts in more detail in Section 6.1. The other components depicted in Figure 3 (i.e., the vault, auction, and oracle contracts) are also re-used from the DSS. We perform two sets of experiments: microbenchmarks for the relay subvariants with varying quorum sizes, and end-to-end experiments for CroCoDai’s main operations: creating new tokens, and cross-chain transfers.
6.1. Smart Contracts